OpenStudy (anonymous):

find the foci of the ellispe with the equation x^2/16+y^2/25=1 a.(+3,0) b.(0,+3) c.(+/41,0) or d. (0,+/41)

OpenStudy (anonymous):

64?

OpenStudy (anonymous):

How'd you get 64?

OpenStudy (anonymous):

i didnt their was a pic and it said 64, and i said 64? as in how is it 64 its not even an option lol

OpenStudy (anonymous):

To find the foci, you use the following equation.. c^2 = a^2 - b^2

OpenStudy (jdoe0001):

so you have $$ \cfrac{x^2}{16}+\cfrac{y^2}{25}=1 \implies \cfrac{(x-0)^2}{4^2}+\cfrac{(y-0)^2}{5^2}=1\\ \text{the distance from the center to any of the focus is "c"}\\ \color{blue}{c= \sqrt{a^2-b^2}} $$ notice your center is (0, 0)

OpenStudy (anonymous):

?

OpenStudy (jdoe0001):

so, what would you get for the "c" distance?

OpenStudy (anonymous):

i have no idea how to do this stuff thats why i came here for the answer, none of my friend know how either

OpenStudy (jdoe0001):

well, we're under the assumption you've covered this material already, otherwise then this wouldn't apply to you

OpenStudy (jdoe0001):

http://www.youtube.com/watch?v=lvAYFUIEpFI

OpenStudy (anonymous):

c^2 = 25 - 16 c^2 = 9 c = 3

OpenStudy (anonymous):

So eliminate c and d

OpenStudy (anonymous):

thank u

OpenStudy (anonymous):

the alebra class i was in we only did have the book and next year we do the next half so i havent got to this yet..